# Joint symbolic aggregate approximation of time series

### 30 Dec 2023 v1

@@ 3.2. Digitization
The next step is referred to as digitization, which we further transformed the resulting polygonal chain $\widetilde{T}$ into symbolic representation (1). This is obtained by lossy compression technique, i.e., vector quantization (VQ), which is often achieved by mean-based clustering. The concept of vector quantization can be referenced in (Gray, 1984; Dasgupta and Freund, 2009). Given an input of $N$ vectors $P=[p_{1},\ldots,p_{N}]\in\mathrm{R}^{\ell\times N}$, VQ seeks a codebook of $k$ vectors, i.e., $C=[c_{1},\ldots,c_{k}]\in\mathrm{R}^{\ell\times k}$ where $k$ is much smaller than $N$. A quality codebook enables sum of squared errors SSE is small enough to an optimal level. Suppose $k$ clusters $S_{1},S_{2},\ldots,S_{k}\subseteq P$ are computed, VQ aims to minimize
@@ 6.2. Multithreading simulation
We used Gaussian noises as the time series for benchmarking. The data generated for the test are of length 100,000 with zero mean and unit standard deviation. We first ran fABBA with $\text{{tol}}=0.01$ and $\alpha=0.05$ to compute the number of symbols it used. This simulation used 358 symbols accordingly. Second, we ran ABBA by feeding the same number of symbols fABBA used to $k$, i.e., $358$ symbols. After that, we run the JABBA (VQ) and JABBA (GA) with varying partitions by the same tol and specifying a consistent hyperparameter setting for digitization, i.e., $k=358$ and $\alpha=0.05$, respectively. The number of threads scheduled for JABBA is set the same as the partition number. The compression rate $\tau_{c}$ computed for all competing algorithms is 0.997, which means that this temporal data are highly compressed in this simulation.

### 02 Jan 2024 v2

@@ 3.2. Digitization
The next step is referred to as digitization, which we further transformed the resulting polygonal chain $\widetilde{T}$ into the symbolic representation in the form of (1).
Following (Elsworth and Güttel, 2020a), prior to digitizing, the tuple lengths and increments are separately normalized by their standard deviations $\sigma_{\texttt{len}}$ and $\sigma_{\texttt{inc}}$, respectively. After that, further scaling is employed by using a parameter scl to assign different weights to the length of each piece $p_{i}$, which denotes importance assigned to its length value in relation to its increment value. Hence, the clustering is effectively performed on the scaled tuples
(4) $p_{1}=\left(\texttt{scl}\frac{\texttt{len}_{1}}{\sigma_{\texttt{len}}},\frac{\texttt{inc}_{1}}{\sigma_{\texttt{inc}}}\right),\ldots,p_{n}=\left(\texttt{scl}\frac{\texttt{len}_{n}}{\sigma_{\texttt{len}}},\frac{\texttt{inc}_{n}}{\sigma_{\texttt{inc}}}\right).$
In particular, if $\texttt{scl}=0$, then clustering will be only performed on the increment values of $P$, while if $\texttt{scl}=1$, the lengths and increments are clustered with equal importance.
The steps after normalization proceed with a lossy compression technique, e.g., vector quantization (VQ), which is often achieved by mean-based clustering. The concept of vector quantization can be referenced in (Gray, 1984; Dasgupta and Freund, 2009). Given an input of $N$ vectors $P=[p_{1},\ldots,p_{N}]\in\mathrm{R}^{\ell\times N}$, VQ seeks a codebook of $k$ vectors, i.e., $C=[c_{1},\ldots,c_{k}]\in\mathrm{R}^{\ell\times k}$ such that $k$ is much smaller than $N$ where each $c_{i}$ is associated with a unique cluster $S_{i}$. A quality codebook enables the sum of squared errors SSE to be small enough to an optimal level. Suppose $k$ clusters $S_{1},S_{2},\ldots,S_{k}\subseteq P$ are computed, VQ aims to minimize
@@ 6.2. Multithreading simulation
We used Gaussian noises as the time series for benchmarking. The data generated for the test are of length 100,000 with zero mean and unit standard deviation. We first ran fABBA with $\text{{tol}}=0.01$ and $\alpha=0.05$ to compute the number of symbols it used. This simulation used 358 symbols accordingly. Second, we ran ABBA by feeding the same number of symbols fABBA used to $k$, i.e., $358$ symbols. After that, we run the JABBA (VQ) and JABBA (GA) with varying partitions by the same tol and specifying a consistent hyperparameter setting for digitization, i.e., $k=358$ and $\alpha=0.05$, respectively. The number of threads scheduled for JABBA is set the same as the partition number. The result shows the compression rate $\tau_{c}$ computed for ABBA, fABBA, JABBA (GA), and JABBA (VQ) are , respectively.